### 16 Mar 2026 v1

# A Kähler and quaternion-Kähler spacetime structure

Abstract
When real Lorentzian spacetime is embedded into a manifold parametrized by higher division algebras (complex or quaternion with Hermitean metric) and the representation constraints of their symmetry groups are made compatible, a set of quantum numbers is generated that is evocative of those of the standard model of particle physics. This is taken here as a hint that in spacetime there is a pseudo-Kähler or pseudo-quaternion-Kähler structure, real spacetime being a submanifold that inherits the symmetry contraints of the larger ambient manifold.
1 Introduction
As stated in the abstract, this note describes a Kähler (KM) and quaternion-Kähler manifold (QKM) perspective for the structure of spacetime. This, of course, might be regarded as a pure mathematical exercise of academic interest. However, there are some hints suggesting that the embedding of real Lorentzian spacetime into a manifold parametrized by higher division algebra $\left(\mathbb{A-}\right)$ numbers (complex or quaternion) might provide, through symmetry considerations, a rationale for the profusion of quantum numbers in the standard model of particle physics.
There are essentially two distinct ways to embed real Lorentzian spacetime $\mathcal{M}_{\mathbb{R}}$ in a higher division algebra manifold $\mathcal{M}_{\mathbb{A}}$. With the reference flat metric $G=\left(1,-1,-1,-1\right)$ one may consider either
$\Lambda^{T}G\Lambda=G$
or
$\Lambda^{\dagger}G\Lambda=G,$
with $\Lambda_{\alpha\beta}\in\mathbb{A}$, for the extended Lorentz group. In the case $\mathcal{\mathbb{A}=}\mathbb{C}$ one obtains in the first case $\Lambda\in SL\left(2,\mathbb{C}\right)\times SL\left(2,\mathbb{C}\right)$, a $12$ parameter group and in the second $\Lambda\in U\left(1,3\right)$, a $16$ parameter group.
The first type of complex extension is the one that has been used in the analytical continuations of the $S$-matrix, interpretation of complex angular momentum and in most studies of complex gravity. On the other hand, the second (Hermitean metric) case is the one that insures a uniform reduction to the usual Lorentz group in each one of the fibers of the Grassmannian of four dimensional frames in $\mathcal{M}_{\mathbb{A}}$.
The two different extensions of the Lorentz group, complemented with translations, provide $\mathbb{A-}$extensions of the Poincaré group with very different topological properties in $\mathcal{M}_{\mathbb{A}}$. A very important difference is the role that spinors play. In the first case spinors in $SL\left(2,\mathbb{C}\right)$ and $SU\left(2,2\right)$ [1] play a central role in the formulation of the global properties of $\mathcal{M}_{\mathbb{A}}$ and on the fibration $\mathbb{C}^{4}\rightarrow\mathcal{M}_{\mathbb{R}}$. In the second (Hermitean) case, spinors exist only as elementary representation states in each four-dimensional frame, but not as global states in $\mathcal{M}_{\mathbb{A}}$, reflecting the fact that $U\left(1,3\right)/SO\left(1,3\right)$ is not a spin manifold. Taking this second point of view, the structure of the representation states of the semidirect product of $U\left(1,3,\mathbb{A}\right)$ with the translations $T\left(4,\mathbb{A}\right)$ has been studied in detail in [2] for $\mathbb{A=C},\mathbb{H},\mathbb{O}$  [^1] In the octonion case the construction of the representations should be interpreted in a quasi-algebra framework[3].. The main conclusions are:
1. There are no half-integer spin elementary states for the extended Poincaré groups;
2. There are integer spin irreducible representations of the larger groups. However, when the symmetry groups are still extended to include the discrete transformations (parity and time inversion), together with positivity of energy in the real Lorentz submanifolds imply the existence of a superselection rule between the integer spin states of the larger groups and those associated to the real Poincaré groups in the real submanifolds.
The conclusion is that spinor states in each real submanifold cannot be ”rotated” in the larger manifold $\mathcal{M}_{\mathbb{A}}$ away from its submanifold $\mathcal{M}_{\mathbb{R}}$. They would be confined to the real submanifolds. On the other hand the integer spin states associated to each real submanifold can only interact with the global integer spin states by $T$-violating interactions.
In the past, the fact that the pseudo-unitary algebras do not possess elementary spinor representations has led some authors to dismiss them as spectrum-generating algebras. However the non-existence of spinor states, as linear irreducible representations, is their most interesting feature. The nonexistence of spinors as linear representations, does not mean that they cannot be implemented in a non-linear way. This may be done by expressing the groups as principal bundles, with structure groups the subgroups of the real submanifolds, and then constructing the associated vector bundles. This has been carried out in [4] for massive and massless states. For example, for massive states the isotropy group is $U\left(3,\mathbb{C}\right)$ for the larger group and $SO\left(3,\mathbb{R}\right)$ for the real submanifold. $U\left(3,\mathbb{C}\right)$ is looked at as a principal bundle with base $W=U\left(3,\mathbb{C}\right)/SO\left(3,\mathbb{R}\right)$ and fiber $SO\left(3,\mathbb{R}\right)$. $W$ is not a spin manifold and to convert it into a spin manifold one has to add to it, by a Whitney sum, an additional $Sp\left(1\right)\simeq SU\left(2\right)$ manifold. In the process new quantum numbers are added to the spinors of the associated vector bundle. In addition one obtains a degeneracy of states with identical quantum numbers in the base $W$.
A simple reasoning, that applies whether or not the base of the principal bundles is a spin manifold, is to require the representations of both the larger group and the groups of the submanifolds to match each other. Therefore, when spinors are implemented in the fibers, some additional spinor quantum numbers have to be added to obtain, by coupling, the integer spin states of the larger group.
In conclusion: requiring in this way the states in the real submanifold to inherit the symmetries of the larger complex group, a set of additional quantum numbers was obtained that is evocative of those of one-generation standard model. As stated in [4] the structure is not exactly the same as the algebra of the standard model. In particular the color-like degenerate states are not linear representations of $SU\left(3\right)$, but independent directions in the base manifold $W.$
A similar discussion has been carried out for the quaternion case in [5]. The quantum number structure that is obtained is evocative of a 33-generations standard model.
Is the quantum number structure of the standard model really a consequence of the symmetries of a larger space where our real spacetime is embedded? Is this an actual embedding or a mathematical fiction? Can it ever be decided? Strangely, so far the best hint comes from the algebraic structure of standard model.
So far all the symmetry considerations, discussed before, concerned the tangent space of spacetime. If one wants to further explore the consequences of embedding real spacetime in larger division algebra spaces, and in particular what consequences it might have for cosmological models, further considerations must be called into play. This will be attempted in the next section for the complex case and in section $3$ for the quaternion case.
2 Universe as a local pseudo-Kähler manifold
The simpler complex embedding of a four-dimensional Lorentzian spacetime (with symmetry group $SO\left(1,3\right)\circledS T\left(4,\mathbb{R}\right)$) is into a flat manifold $\mathcal{M}_{\mathbb{C}}^{0}$ with symmetry group $U\left(1,3\right)\circledS T\left(4,\mathbb{C}\right)$ with coordinates
$z^{\mu}=x^{\mu}+iy^{\mu}$ (1)
and Hermitean metric
$g=\eta_{\overline{\mu}\nu}d\overline{z}^{\mu}dz^{\nu},$ (2)
$\left\{X^{\mu}=\frac{\partial}{\partial x^{\mu}};Y^{\nu}=\frac{\partial}{\partial y^{\nu}}\right\}$ being a frame for the tangent bundle $T\mathcal{M}_{\mathbb{C}}^{0}$, with multiplication by $i$ a $J$ endomorphism (complex structure)
$JX^{\mu}=Y^{\mu},\hskip 28.45274ptJY^{\mu}=-X^{\mu}.$
The metric satisfies
$g\left(JX,JY\right)=g\left(X,Y\right)$
$X,Y\in T\mathcal{M}_{\mathbb{C}}^{0}$ and the antisymmetric form
$\Omega(X,Y)=g(X,JY)$
is closed, because $J$ is parallel ($\nabla J=0$). Hence $\mathcal{M}_{\mathbb{C}}^{0}$ is a pseudo-Kähler manifold.
All symmetry considerations, here and in Ref.[4], assumed real spacetime as embedded in a larger manifold with local symmetry group the semidirect product of $U\left(1,3,\mathbb{C}\right)$ and $T\left(4,\mathbb{C}\right)$, the abelian four complex spacetime translations. This is the complex Poincaré group [^2] Also called, by some authors [6], the complex Poincaré group with real metric. Hermitean metric would be better, consistent with the Kähler embedding.. However, like in many other cases [7] [8], it is unlikely that this unstable algebra be an actual symmetry of Nature. The commutation relations of the Lie algebra of $U\left(1,3,\mathbb{C}\right)\circledS T\left(4,\mathbb{C}\right)$ are
$\left[M_{\mu\nu},M_{\rho\sigma}\right]$ $=$ $-M_{\mu\sigma}\eta_{\nu\rho}-M_{\nu\rho}\eta_{\mu\sigma}+M_{\nu\sigma}\eta_{\rho\mu}+M_{\mu\rho}\eta_{\nu\sigma}$
$\left[M_{\mu\nu},N_{\rho\sigma}\right]$ $=$ $-N_{\mu\sigma}\eta_{\nu\rho}+N_{\nu\rho}\eta_{\mu\sigma}+N_{\nu\sigma}\eta_{\rho\mu}-N_{\mu\rho}\eta_{\nu\sigma}$
$\left[N_{\mu\nu},N_{\rho\sigma}\right]$ $=$ $M_{\mu\sigma}\eta_{\nu\rho}+M_{\nu\rho}\eta_{\mu\sigma}+M_{\nu\sigma}\eta_{\rho\mu}+M_{\mu\rho}\eta_{\nu\sigma}$
$\left[M_{\mu\nu},K_{\rho}\right]$ $=$ $-\eta_{\nu\rho}K_{\mu}+\eta_{\mu\rho}K_{\nu}$
$\left[M_{\mu\nu},H_{\rho}\right]$ $=$ $-\eta_{\nu\rho}H_{\mu}+\eta_{\mu\rho}H_{\nu}$
$\left[N_{\mu\nu},K_{\rho}\right]$ $=$ $-\eta_{\nu\rho}H_{\mu}-\eta_{\mu\rho}H_{\nu}$
$\left[N_{\mu\nu},H_{\rho}\right]$ $=$ $\eta_{\nu\rho}K_{\mu}+\eta_{\mu\rho}K_{\nu}$
$\left[K_{\mu},H_{\rho}\right]$ $=$ $0$ (3)
with $\mu,\nu,\rho,\sigma\in\left\{0,1,2,3\right\}$, the antisymmetric $M_{\mu\nu}$ are the generators of real rotations and boosts in $4$ dimensions with metric $\eta=(1,-1,-1,-1)$, the symmetric $N_{\mu\nu}$ the generators of complex transformations and $H_{\mu}$ and $K_{\mu}$ the generators of real and complex translations.
The minimal deformation that stabilizes this algebra is [^3] Reduction (contraction) to the complex Poincaré group would be obtained by $\frac{M_{\mu 4}}{R}\rightarrow P\mu$, when $R\rightarrow\infty$
$\left[M_{ab},M_{cd}\right]$ $=$ $-M_{ad}\eta_{5bc}-M_{bc}\eta_{5ad}+M_{bd}\eta_{5ca}+M_{ac}\eta_{5bd}$
$\left[M_{ab},N_{cd}\right]$ $=$ $-N_{ad}\eta_{5bc}+N_{bc}\eta_{5ad}+N_{bd}\eta_{5ca}-N_{ac}\eta_{5bd}$
$\left[N_{ab},N_{cd}\right]$ $=$ $M_{ad}\eta_{5bc}+M_{bc}\eta_{5ad}+M_{bd}\eta_{5ca}+M_{ac}\eta_{5bd}$ (4)
with $a,b,c,d\in\left\{0,1,2,3,4\right\}$ and metric $\eta_{5}=(1,-1,-1,-1,g_{44})$. With $g_{44}$ either 1-1 or +1+1 these are the stable algebras of $U\left(1,4,\mathbb{C}\right)$ or $U\left(2,3,\mathbb{C}\right)$. They are the symmetry algebras of flat complex pseudo-Euclidean spaces in $1+4$ or $2+3$ dimensions.
To obtain a space $\mathcal{M}_{\mathbb{C}^{4}}$ of four complex dimensions with the same symmetry, one starts from a complexified flat five dimensional space $\mathcal{M}_{\mathbb{C}^{5}}$ with metric $\eta_{5}$ which, by the same considerations as above, is a pseudo-Kähler space. Then, spaces of smaller complex dimension with the same symmetry are obtained in submanifolds defined by invariant homogeneous polynomials. A simple choice would be to restrict the coordinates by
$\left|z_{0}\right|^{2}-\left|z_{1}\right|^{2}-\left|z_{2}\right|^{2}-\left|z_{3}\right|^{2}+g_{44}\left|z_{4}\right|^{2}=-R^{2}.$ (5)
Depending on the $\mp$ sign of $g_{44}$, the four-dimensional spacetimes $\mathcal{M}_{\mathbb{C}^{4}}$ that are obtained are complex versions of de Sitter or anti-de Sitter spaces.
Let, for definiteness, $g_{44}=-1$ and parametrize the coordinates by
$z^{0}$ $=$ $R\sinh\tau e^{i\psi_{0}}$
$z^{1}$ $=$ $R\cosh\tau\cos\theta_{1}e^{i\psi_{1}}$
$z^{2}$ $=$ $R\cosh\tau\sin\theta_{1}\cos\theta_{2}e^{i\psi_{2}}$
$z^{3}$ $=$ $R\cosh\tau\sin\theta_{1}\sin\theta_{2}\cos\theta_{3}e^{i\psi_{3}}$
$z^{4}$ $=$ $R\cosh\tau\sin\theta_{1}\sin\theta_{2}\sin\theta_{3}e^{i\psi_{3}}.$ (6)
With this choice (5) is satisfied and, from the flat invariant length $ds_{5}^{2}=d\overline{z_{0}}dz_{0}-d\overline{z_{1}}dz_{1}-d\overline{z_{2}}dz_{2}-d\overline{z_{3}}dz_{3}-d\overline{z_{4}}dz_{4}$ for the five-dimensional $\eta_{5}$ metric, one obtains for the complex four dimensional submanifold $\mathcal{M}_{\mathbb{C}^{4}}$
$\frac{ds_{4}^{2}}{R^{2}}$ $=$ $d\tau^{2}+\sinh^{2}\tau d\psi_{0}^{2}-\cosh^{2}\tau\left(d\theta_{1}^{2}+\sin^{2}\theta_{1}d\theta_{2}^{2}+\sin^{2}\theta_{1}\sin^{2}\theta_{2}d\theta_{3}^{2}\right$. (7)
$\left.+\cos^{2}\theta_{1}d\psi_{1}^{2}+\sin^{2}\theta_{1}\cos^{2}\theta_{2}d\psi_{2}^{2}+\sin^{2}\theta_{1}\sin^{2}\theta_{2}d\psi_{3}^{2}\right),$
the factor that multiplies $\cosh^{2}\tau$ being the measure of a complex $3$-sphere, $d\Omega_{3\mathbb{C}}$. In these coordinates the complex structure $J$ is
$\tau\overset{J}{\rightarrow}\tau$
$\psi_{i}\overset{J}{\rightarrow}\psi_{i}+\pi/2$
and $J\left(T_{z}\mathcal{M}_{\mathbb{C}^{4}}\right)=T_{z}\mathcal{M}_{\mathbb{C}^{4}}$ for all $z\in\mathcal{M}_{4}$. Therefore $\mathcal{M}_{\mathbb{C}^{4}}$, being a complex submanifold of $\mathcal{M}_{\mathbb{C}^{5}}$, is also a pseudo-Kähler manifold with respect to its induced structure [9].
$\mathcal{M}_{\mathbb{C}^{4}}$ is a two-time field which, with the change
$\cos T=\frac{1}{\cosh\tau};\hskip 28.45274pt-\frac{\pi}{2}<T<\frac{\pi}{2},$
$ds_{4}^{2}=\frac{R^{2}}{\cos^{2}\tau}\left(dT^{2}+\sin^{2}Td\psi_{0}^{2}-d\Omega_{3\mathbb{C}}\right)$
may be (conformally) represented by a three-dimensional (Penrose-like) diagram (Fig.1), each horizontal plane being a complex 33-sphere and each longitudinal line a complex 22-sphere.
Figure 1: Penrose-like diagram for complex de Sitter spacetime. Horizontal planes (red) are complex 3-spheres and longitudinal lines (blue) complex 2-spheres. Light rays propagate on planes at 45 degrees
In a real Lorentzian submanifold $\left\{\psi_{\mu}=0\right\}$, this $\mathcal{M}_{\mathbb{C}^{4}}\mathcal{-}$ structure implies a cosmological constant $\Lambda=\frac{3}{R^{2}}$. In the present symmetry context, a finite $R$ appears has a natural requirement of stability of the overall kinematical algebra. A value $R=\infty$ $\left(\Lambda=0\right)$ would correspond to an improbable exceptional choice of an unstable algebra. Hence, to look for a mysterious dark energy as the origin of a non-zero $\Lambda$ makes as must sense as to look for the reasons for the specific values of the speed of light or of the Planck constant. We do not know why these constants have the values they have, we only know that we live in a generic Universe, rather than an exceptional one with the physical constants adjusted to unstable algebras.
The strictly local symmetries are well approximated by the semidirect product $U\left(1,3,\mathbb{C}\right)\circledS T\left(4,\mathbb{C}\right)$ and stability (genericity) of the symmetry requires $U\left(1,4,\mathbb{C}\right)$. However, as stated in the title of this section, this might also be only a local structure, other constraints might operate at larger scales.
Complexified forms of spacetime with or without hermitean metrics have been discussed before by many authors , going back to Einstein [10]. They appeared either as extensions of the space of solutions or completeness of geodesics and string propagation, non-local interpretations of Minkowski or de Sitter space (the twistor approach), etc ([11]-[21]). What apparently had not been realized before was the importance of the distinct nature of the representation spaces for the pseudo-orthogonal and the pseudo-unitary groups, when they operate as symmetry groups of the real and the complex spacetimes. And also that when the symmetries of the complex space are forced upon the real representation states, new quantum numbers emerge from the bundle Whitney sums.
The non-existence of spinors as representation states when one goes from the real to the (Hermitean ) complex or quaternionic spacetime has a simple topological reason. Spinor representations are associated to the spin groups, which are double covers of orthogonal or pseudo-orthogonal groups, $Spin(N)$ being a double cover of $SO(N)$ that allows projective representations of $SO(N)$ to be lifted to linear representations of $Spin(N)$. In contrast, $U(N)$ is simply connected and has no nontrivial double cover that would give rise to spinor representations.
The complexified de Sitter manifold $\mathcal{M}_{\mathbb{C}^{4}}$ would be a complex universe without matter, the constant $R$ being simply a constant of Nature revealing the universe as generic, not an exceptional one ($R=\infty$). In this sense it is independent of the matter content of the Universe. However, to construct cosmological models the presence of matter should be taken into account. Here also, the representation state constraints, previously obtained [2], play a role. Namely, bulk matter, that is, matter that can be associated to the complex universe as a whole, is only bosonic matter, fermionic matter being confined to the Lorentzian four-dimensional submanifolds. In addition, the complex bulk matter will only interact with the submanifolds’ matter (either bosonic or fermionic) by $T$-violating interactions. This has implications on the dynamics of the complex space because singular terms (delta-restricted to the submanifolds) must be represented in the stress-energy tensor. On the other hand the dynamics on the submanifolds should also take into account the $T$-violating interactions arising from an eventual distribution of bulk matter [22]. This $T$-violating interactions of the bulk bosonic matter, with the matter associated to the submanifold representations, might be interpreted as a gravitational interaction, or might be something different. A simple reasoning, based on the Birkoff theorem implies that, at large distances from the center, the rotation velocity of the star galaxies may actually grow with the distance. This comes from the fact that bulk matter is spread over a six dimensional spatial volume.
A real four-dimensional Lorentzian spacetime is trivially embedded, as a submanifold, in $\mathcal{M}_{\mathbb{C}^{4}}$ by $\left\{\psi_{\mu}=0\right\}$. A relevant question is what other types of submanifolds can be identified. Indeed, in a Kähler manifold there are several typical interesting families of submanifolds, namely complex, totally real, CR and slant submanifolds. They are classified by the action of the complex structure on the tangent bundle. In complex submanifolds $\mathcal{M}$ the tangent bundle is invariant under the complex structure $J$ ($J\left(T_{z}\mathcal{M}\right)=T_{z}\mathcal{M}$). This is the case for $\mathcal{M}_{\mathbb{C}^{4}}$ as a submanifold of $\mathcal{M}_{\mathbb{C}^{5}}$. In totally real manifolds $J\left(T_{z}\mathcal{M}\right)\subset T_{z}^{\bot}\mathcal{M}$, that is, $J$ maps tangent spaces into the normal bundle. A submanifold $\mathcal{N}$ is called a CR-submanifold if there exist a holomorphic distribution $H=\left\{H_{z}=T_{z}\mathcal{N}\cap J(T_{z}\mathcal{N}):z\in\mathcal{N}\right\}$ and a totally real distribution $H^{\perp}$ such that $T\mathcal{N}=H\oplus H^{\perp}$.
Real submanifolds inherit important differential geometry structures from the Kähler manifold [23]. With local metric $\eta=(1,-1,-1,-1)$ dimension counting implies that real four-dimensional submanifolds will necessarily be Lorentzian manifolds. However other types of four-dimensional submanifolds are possible in the complexified de Sitter spacetime, four-dimensional universes without time or with two time-like directions.
The extension of spacetime coordinates to higher division algebras [2] and the non-linear inheritance of the symmetries of the larger spaces [4] [5] being at the origin of the quantum numbers of the standard model, would imply that real space time is embedded in a higher dimensional space, eight-dimensional in the complex extension case and sixteen-dimensional in the quaternionic case. Also, the best understood string theory models involve ten-dimensional spacetimes, with the extra dimensions being curled up on small compact manifolds. Here instead, no compactification of the extra dimensions is involved, the relative independence of the four-dimensional submanifolds being a representation theory effect of the symmetry groups. There are however possible $T$-violating interactions with some form of bulk bosonic matter of the extended space.
3 The quaternion-Kähler structure
In a quaternion-Kähler manifold [24] one has three anticommuting complex structures. A quaternion-Kähler manifold may or may not be a Kähler manifold, it all depends on the existence of a global complex structure.
Here the four-dimensional Lorentzian spacetime would be embedded into a quaternion manifold with strictly local symmetry $U\left(1,3,\mathbb{H}\right)\circledS T\left(4,\mathbb{H}\right)$ and real dimension $16$. The restrictions on representation spaces and interactions are similar to the complex case [2]. Stability of the algebras, that is, genericity of the symmetries, leads to $U\left(1,4,\mathbb{H}\right)$ which, with a construction similar to the complex case implies a quaternion-Kähler version of de Sitter space. For the coordinates one may use the same form as in (6) with $e^{i\psi_{\mu}}$ replaced by unit quaternions.
Could this embedding be physically relevant? As shown in [5] this embedding generates a representation manifold of elementary states with a structure compatible with the three generations of the standard model. However the higher dynamical masses of the second and third generation, suggest that effects of the quaternionic embedding might only be relevant for higher energies and not, for example, for the galaxy rotation problem. And could these embeddings be further extended to higher dimensions? The octonionic case was also discussed in [2], with the non-associative octonions interpreted in the quasi-algebra sense [3]. It might suggest a still deeper sub-quark structure.